OpenStudy (2yoututu11):

What is the area of the rectangle? A. 25.5 square feet B. 28 square feet C. 48 square feet D. 51 square feet

OpenStudy (albert0898):

Area of rectangle = l * w Area = 8.5 * 6 Area = 51